OpenStudy (anonymous):

Two negative integers have a sum of -7 and a product of 6. What are the integers?

OpenStudy (anonymous):

-1, -6

OpenStudy (anonymous):

(-1) + (-6) = -7 -1 x -6 = 6

OpenStudy (anonymous):

Thank you (: